OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs of 20 cm and 21 cm. 41 cm 841 cm 6 cm 29 cm

OpenStudy (anonymous):

\[H= \sqrt{(20)^2+(21)^2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[c^2=a^2+b^2\]

OpenStudy (anonymous):

so i have to find out what H is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[H= \sqrt{(20^2)+(21^2)}\]

OpenStudy (anonymous):

20 * 20 = 400 21 * 21 = 441 add those 400 + 441 then take the sqrt and that's the missing angle

OpenStudy (anonymous):

841cm

OpenStudy (anonymous):

\[\sqrt{841cm}\] sqrt it and the answer is?

OpenStudy (anonymous):

so then 29cm

OpenStudy (anonymous):

yes ma'am